# Lost in the FoG: Pitfalls of Models for Large-Scale Hydrogen Distributions

### v4

@@ 5.3.2 Predicted impact on recent observations
However, the ${\rm H}$i distribution in IllustrisTNG would not be perfectly measured by 21cm intensity mapping experiments due to re-gridding (Cunnington & Wolz, 2024) and the beams of the instruments themselves. In previous sections, we remained agnostic to instrument-specific effects in order to keep our results general and to focus on errors originating from models. In Fig. 10, however, we examine the impact of model errors on the $\Omega_{\mathrm{HI}}$ constraints from specific observations. To imitate how each instrument would observe IllustrisTNG’s ${\rm H}$i distribution, we apply a Gaussian filter to the simulated redshift-space ${\rm H}$i density grid, following Wolz et al. (2017):
@@
Here, $d_{c}$, $c$, $N_{\rm{grid}}$, and $\nu$ are the co-moving distance, speed of light, number of grid points along one axis, and redshifted 21cm line, respectively. $D_{\rm{dish}}$ represents the diameter of the dish for each radio telescope — we use 100m for Greenbank Telescope (GBT, Masui et al., 2013; Switzer et al., 2013) and 13.5m for MeerKAT (MKT, Wang et al., 2021a). We assume a constant beam size along the line of sight, as the box is at one redshift, and apply the smoothing on each two-dimensional slice perpendicular to the line-of-sight ($z$) independently.
We also adjust our model to accommodate the beam, following Ando et al. (2019):
@@
Overall, the revisions improve the agreement between the $\Omega_{\mathrm{HI}}$ constraints from ${\rm H}$i-galaxy cross-power spectra and other techniques (Zwaan et al., 2005; Rao et al., 2006; Lah et al., 2007; Martin et al., 2010; Braun, 2012; Rhee et al., 2013; Hoppmann et al., 2015; Rao et al., 2017; Jones et al., 2018; Bera et al., 2019; Hu et al., 2019; Chowdhury et al., 2020). This is particularly true for the constraints from W22, while revised $\Omega_{\mathrm{HI}}$ values from C22, CHIME23, and C24 are still consistent with the other constraints. The $\Omega_{\mathrm{HI}}$ constraint from CHIME $\times$ ELG is revised downward, although it is not shown in Fig. 10 due to its unusually large value — CHIME23 attribute this to a chance fluctuation and the prior volume effect (see their section 6.3).
@@ 6.1 What is the impact of simulation technique on the hydrogen distribution?
This work Hydro - 5.9 ×$\times 10^{7}$ 0.07 - 1.28
Ando et al. (2019) Hydro - 2.9 ×$\times 10^{7}$ 0.26 - 1.22
VN18 Hydro - 7.5 ×$\times 10^{6}$ 0.15 - 1.49
Wolz et al. (2016) N-body SAM $1.2\times 10^{9}$ $0.1$ $M_{h}\geq 10^{10}$ $\sim 1.8$
Spinelli et al. (2020) N-body SAM $1.2\times 10^{9}$ $0.015$ $M_{h}\gtrsim 10^{10},M_{\star}\gtrsim 10^{8}$ 1.22
Wang et al. (2021b) N-body SAM $5\times 10^{8}$ 0.05 $M_{h}\geq 1.5\times 10^{10}$ 1.26
Guo et al. (2023) N-body SAM $2\times 10^{8}$ 0.05 $M_{\mathrm{HI}}\geq 10^{6.9}$ 1.10
Spinelli et al. (2020) N-body SAM $10^{7}$ $0.1$ $M_{h}\gtrsim 10^{8},M_{\star}\gtrsim 10^{6}$ 1.31
Wang et al. (2021b) N-body HIHM $5\times 10^{8}$ 0.05 $M_{h}\geq 1.5\times 10^{10}$ 1.48
Sarkar et al. (2016) N-body HIHM $10^{8}$ 0.065 - $\sim 0.9$
VN18 N-body HIHM 7.5 ×$\times 10^{6}$ 0.15 - $\sim 1.2$
Pénin et al. (2018) Analytical - - 0.01 - $\sim$0.82
Umeh (2017) Analytical - - 0.01 - $\sim$0.86
@@
In Section 5, we found that IllustrisTNG predicts that general large-scale models (Equations 12-13) introduce significant errors in cosmological constraints. However, simulated ${\rm H}$i distributions can be sensitive to the included galaxy formation physics (Park et al., 2014; Li et al., 2024; Marasco et al., 2025). This raises a key question: does IllustrisTNG’s galaxy formation model inherently yield complicated ${\rm H}$i distributions? We explore this question by comparing results from some selected computational and theoretical works in Table 5, broadly characterized by their $z=1$ ${\rm H}$i bias.
@@
We examine results from three methodologies: Hydro, N-body, and Analytical. The first two, Hydro and N-body, describe simulations that differ in their handling of baryons. Hydrodynamical simulations like IllustrisTNG include baryonic physics on-the-fly, whereas N-body simulations defer baryon modeling to post-processing, typically through a ${\rm H}$i-halo mass relation (HIHM) or semi-analytic model (SAM). In the remaining methodology, Analytical, perturbation theory (Bernardeau et al., 2002) and a halo model are employed to study ${\rm H}$i distributions without a simulation.
@@
${\rm H}$i bias values in the first methodology, Hydro, appear to agree across different simulations at $z=1$ (also see figure 2 from Ando et al., 2019), although they show notable dependence on resolution. We note that our model tests are robust against these resolution effects since they are dependent on the ${\rm H}$i bias shape, rather than its precise value (Appendix A). We attribute the Hydro resolution dependence to differences in how ${\rm H}$i occupies the largest halos. VN18 found that massive halos in coarser simulations tended to contain less ${\rm H}$i than finer ones, in turn suppressing ${\rm H}$i clustering. This agrees with other works that studied the resolution dependence of star-formation in IllustrisTNG. For example, Donnari et al. (2021) found that the quenching of the massive satellites ($M_{\star}\geq 10^{10}M_{\odot}$) in these massive host halos is enhanced in the coarser IllustrisTNG simulations, in turn suggesting reduced ${\rm H}$i abundance.
@@
SAMs exhibit a similar resolution trend in their ${\rm H}$i clustering, albeit for different reasons. Instead of a reduction of ${\rm H}$i occupation in large halos, SAMs of coarser-resolution simulations possess enhanced ${\rm H}$i occupation in small halos (see figure 6 in Spinelli et al., 2020). In both SAMs and Hydro, ${\rm H}$i occupies smaller halos on average, suppressing its clustering. The ${\rm H}$i bias value also depends on the particular SAM implementation, as demonstrated by Wolz et al. (2016) and Guo et al. (2023). Identifying the particular SAM components that cause these differences is left to future work.
@@
HIHMs, conversely, have a non-trivial dependence on resolution and the employed HIHM. The resolution dependence is clearest when comparing VN18 and Wang et al. (2021b), which adopt the same HIHM but still find ${\rm H}$i biases that differ by $\sim 0.3$. Wang et al. (2021b) reasoned that the discrepancy arises because halos below their resolution threshold contribute non-negligibly to the overall ${\rm H}$i distribution. However, even at similar resolutions, Sarkar et al. (2016) and Wang et al. (2021b) find biases that differ by $\sim 0.5$, implying that the ${\rm H}$i bias is also dependent on the adopted HIHM.
@@
Comparisons among the methodologies indicate that ${\rm H}$i bias values are sensitive to secondary halo properties, such as halo age or assembly history (Gao et al., 2005). Methodologies like Hydro and SAMs that natively incorporate these assembly effects generally converge within $1.51.2<b_{\mathrm{HI}}<$ 1.5, particularly at comparable resolutions (Ando et al., 2019). Marasco et al. (2016) found that SAMs have stronger ${\rm H}$i occupation in dense environments as compared to Hydro, which may influence the remaining differences between Hydro and SAMs. Conversely, methodologies that neglect assembly effects are either substantial outliers from Hydro and SAMs, like Analytical, or vary dramatically between implementations, like HIHM. However, in the case of HIHM, VN18 showed that its overall normalization and the fraction of ${\rm H}$i outside of halos can also contribute to this disparity in behavior.
@@ 6.2 Why is the FoG effect such a large source of error?
One known auxiliary damping source is the neglected nonlinear Kaiser terms in Equations 12-13, which are typically included as corrections to the matter power spectrum (Scoccimarro, 2004; Taruya et al., 2010). Ando et al. (2019) demonstrated that strictly linear models have reduced PVDs relative to models with these terms (see their figure 5). However, these models necessarily assume that nonlinearities manifest the same in ${\rm H}$i and matter (i.e., no ${\rm H}$i velocity bias), and furthermore including these nonlinearities can actually worsen model performance at $0.5\lesssim k\lesssim 1$ h cMpc-1. Given these complications, it is uncertain if the nonlinear contributions to the Kaiser term alone can capture the trends seen in Fig. 11.
Differences in the small- and large-scale behavior of the ${\rm H}$i PVD may also give rise to the behavior in Fig. 11. The ${\rm H}$i PVD can be understood in a halo framework as contributions of three terms, each dominating at different scale regimes: inter-halo (large scales), intra-halo (intermediate), and intra-galactic (small) velocity dispersions (Slosar et al., 2006; Sarkar & Bharadwaj, 2019). The last contribution is unique to ${\rm H}$i intensity maps as compared to point-source tracers like galaxies, which would not have intra-galactic dispersions (P23; Li et al., 2024). Osinga et al. (2024) showed that the ${\rm H}$i power spectrum with and without these intra-galactic dispersions diverge at $k\sim 0.5$ h cMpc-1 in IllustrisTNG, with the latter asymptoting to a constant shot noise floor. This comparison demonstrates that intra-galactic dispersions become important at $k\gtrsim 0.5$ h cMpc-1 and the real-space shot noise term is damped only by intra-galactic dispersions [^4] Strictly speaking, shot noise only manifests on scales where random noise emerges and thus cannot be “damped” by RSDs. A more precise interpretation is that ${\rm H}$i fills a larger fraction of intra-halo spaces in redshift space, effectively lowering the shot noise floor compared to real space. Consequently, the actual redshift-space shot noise floor should resurface at smaller scales.. Given this conclusion, we propose that ${\rm H}$i power spectra be modeled phenomenologically with two separate FoG damping terms:
@@
Here, $P^{\mathrm{SN}}$ symbolizes the shot noise term from the real-space ${\rm H}$i auto power spectrum. $\sigma_{\mathrm{HI}}$ represents the ${\rm H}$i PVD, which would include all three inter-halo, intra-halo, and intra-galactic components, whereas $\sigma_{\mathrm{int}}$ only includes the intra-galactic dispersions. The two distinct damping terms would reproduce the behavior in Fig. 11 if $\sigma_{\mathrm{int}}<\sigma_{\mathrm{m}}<\sigma_{\mathrm{HI}}$, assuming that shot noise is not significant to the matter power spectrum. A thorough understanding of the distribution of the galactic ${\rm H}$i emission profiles can improve the modeling of $\sigma_{\mathrm{int}}$ (Li et al., 2024).
@@ 7 Conclusion
    The error from assuming a constant bias reaches $\gtrsim 10\%$ at $k\sim 0.2-0.3$ h cMpc-1 for ${\rm H}$i and all galaxy populations. The exceptions are the $z=1$ ${\rm H}$i bias and $z=0$ blue galaxy bias, which remain constant at $<10\%$ level until $k\sim 1$ h cMpc-1 because halo occupation effects and nonlinear clustering coincidentally cancel.
@@
    The error from assuming a constant ${\rm H}$i-blue galaxy correlation coefficient reaches $10\%$ at $k\sim 0.3$ h cMpc-1. Conversely, ${\rm H}$i-red deviates from its large-scale value by $>10\%$ on all scales probed by IllustrisTNG ($k\gtrsim 0.05$ h cMpc-1).
@@
However, we only studied the performance of these models using a single suite of simulations. Completing similar studies with other simulations would test how its architecture impacts the resulting tracer distributions and the performance of large-scale ${\rm H}$i models. Furthermore, we only examine how models perform in the monopole — higher-order cross-correlations between ${\rm H}$i and galaxies can provide even more constraining power, but require even more sophisticated models (e.g., Witzemann et al., 2019; Chand et al., 2024).
Acknowledgements
We would like to thank Steven Cunnington and Laura Wolz for their advice, comments, and helpful resources in replicating their work. We also thank Alkistis Pourtsidou and Yi-chao Li for their useful discussions. Research was performed in part using the compute resources and assistance of the UW-Madison Center For High Throughput Computing (CHTC) in the Department of Computer Sciences. The CHTC is supported by UW-Madison, the Advanced Computing Initiative, the Wisconsin Alumni Research Foundation, the Wisconsin Institutes for Discovery, and the National Science Foundation. We also acknowledge the University of Maryland supercomputing resources (http://hpcc.umd.edu) made available for conducting the research reported in this paper. Other software used in this work include: numpy (Harris et al., 2020), matplotlib (Hunter, 2007), seaborn (Waskom, 2021), h5py (Collette, 2013), colossus (Diemer, 2018) and pyfftw (Gomersall, 2021). This research was supported in part by the National Science Foundation under Grant numbers 2206690 and 2338388. BD furthermore acknowledges support from the Sloan Foundation.
@@ Acknowledgements
Data from the figures in this publication are available upon reasonable request to the authors. The IllustrisTNG data is available at www.tng-project.org/data.

### v5

@@ 5.3.2 Predicted impact on recent observations
However, the $\rm H$}i distribution in IllustrisTNG would not be perfectly measured by 21cm intensity mapping experiments due to re-gridding (S. Cunnington & L. Wolz, 2024) and the beams of the instruments themselves. In previous sections, we remained agnostic to instrument-specific effects in order to keep our results general and to focus on errors originating from models. In Fig. 10, however, we examine the impact of model errors on the $\Omega_{\mathrm{HI}}$ constraints from specific observations. To imitate how each instrument would observe IllustrisTNG’s $\rm H$}i distribution, we apply a Gaussian filter to the simulated redshift-space $\rm H$}i density grid, following L. Wolz et al. (2017):
@@
Here, $d_{c}$, $c$, $N_{\rm{grid}}$, and $\nu$ are the co-moving distance, speed of light, number of grid points along one axis, and redshifted 21cm line, respectively. $D_{\rm{dish}}$ represents the diameter of the dish for each radio telescope — we use 100m for Greenbank Telescope (GBT, K. W. Masui et al., 2013; E. R. Switzer et al., 2013) and 13.5m for MeerKAT (MKT, J. Wang et al., 2021). We assume a constant beam size along the line of sight, as the box is at one redshift, and apply the smoothing on each two-dimensional slice perpendicular to the line-of-sight ($z$) independently.
We also adjust our model to accommodate the beam, following R. Ando et al. (2019):
@@
Fig. 10 contains $\Omega_{\mathrm{HI}}$ constraints from three techniques other than $\rm H$}i intensity mapping (black, M. A. Zwaan et al., 2005; S. M. Rao et al., 2006; P. Lah et al., 2007; A. M. Martin et al., 2010; R. Braun, 2012; J. Rhee et al., 2013; L. Hoppmann et al., 2015; S. M. Rao et al., 2017; M. G. Jones et al., 2018; A. Bera et al., 2019; W. Hu et al., 2019; A. Chowdhury et al., 2020). At low redshifts $(z\approx 0)$, the $\rm H$}i content of galaxies is measured directly in blind surveys, obtaining $\Omega_{\mathrm{HI}}$ by integrating the observed $\rm H$}i mass function. Detections of individual galaxies in a blind survey are difficult at $0.1\lesssim z\lesssim 1$, so the $\rm H$}i of optically-selected galaxies is measured via stacking, with $\Omega_{\mathrm{HI}}$ measurements requiring corrections for incompleteness. Finally, at high redshifts ($z\gtrsim 1$), damped Lyman-$\alpha$ absorption yield column densities that can be integrated to produce $\Omega_{\mathrm{HI}}$.
Overall, the revisions in Fig. 10 improve the agreement between the $\Omega_{\mathrm{HI}}$ constraints from intensity mapping and other techniques. This is particularly true for the constraints from W22, while revised $\Omega_{\mathrm{HI}}$ values from C22, CHIME23, and C24 are still consistent with the other constraints. The $\Omega_{\mathrm{HI}}$ constraint from CHIME $\times$ ELG is revised downward, although it is not shown in Fig. 10 due to its unusually large value — CHIME23 attribute this to a chance fluctuation and the prior volume effect (see their section 6.3).
@@ 6.1 What is the impact of simulation technique on the hydrogen distribution?
This work Hydro - 5.9 $\times 10^{7}$ 0.07 - 1.28
R. Ando et al. (2019) Hydro - 2.9 $\times 10^{7}$ 0.26 - 1.22
VN18 Hydro - 7.5 $\times 10^{6}$ 0.15 - 1.49
L. Wolz et al. (2016) N-body SAM $1.2\times 10^{9}$ $0.1$ $M_{h}\geq 10^{10}$ $\sim 1.8$
M. Spinelli et al. (2020) N-body SAM $1.2\times 10^{9}$ $0.015$ $M_{h}\gtrsim 10^{10},M_{\star}\gtrsim 10^{8}$ 1.22
Z. Wang et al. (2021) N-body SAM $5\times 10^{8}$ 0.05 $M_{h}\geq 1.5\times 10^{10}$ 1.26
H. Guo et al. (2023) N-body SAM $2\times 10^{8}$ 0.05 $M_{\mathrm{HI}}\geq 10^{6.9}$ 1.10
M. Spinelli et al. (2020) N-body SAM $^{7} 0.10.1$ $M_{h}\gtrsim 10^{8},M_{\star}\gtrsim 10^{6}$ 1.31
Z. Wang et al. (2021) N-body HIHM $5\times 10^{8}$ 0.05 $M_{h}\geq 1.5\times 10^{10}$ 1.48
D. Sarkar et al. (2016) N-body HIHM $10^{8}$ 0.065 - $\sim 0.9$
VN18 N-body HIHM 7.5 $\times 10^{6}$ 0.15 - $\sim 1.2$
A. Pénin et al. (2018) Analytical - - 0.01 - $\sim$0.82
O. Umeh (2017) Analytical - - 0.01 - $\sim$0.86
@@
In Section 5, we found that IllustrisTNG predicts that general large-scale models (Equations 12-13) introduce significant errors in cosmological constraints. However, simulated $\rm H$}i distributions can be sensitive to the included galaxy formation physics (J. Park et al., 2014; X. Li et al., 2024; A. Marasco et al., 2025). This raises a key question: does IllustrisTNG’s galaxy formation model inherently yield complicated $\rm H$}i distributions? We explore this question by comparing results from some selected computational and theoretical works in Table 5, broadly characterized by their $z=1$ $\rm H$}i bias.
@@
We examine results from three methodologies: Hydro, N-body, and Analytical. The first two, Hydro and N-body, describe simulations that differ in their handling of baryons. Hydrodynamical simulations like IllustrisTNG include baryonic physics on-the-fly, whereas N-body simulations defer baryon modeling to post-processing, typically through a $\rm H$}i-halo mass relation (HIHM) or semi-analytic model (SAM). In the remaining methodology, Analytical, perturbation theory (F. Bernardeau et al., 2002) and a halo model are employed to study $\rm H$}i distributions without a simulation.
@@
$\rm H$}i bias values in the first methodology, Hydro, appear to agree across different simulations at $z=1$ (also see figure 2 from R. Ando et al., 2019), although they show notable dependence on resolution. We note that our model tests are robust against these resolution effects since they are dependent on the $\rm H$}i bias shape, rather than its precise value (Appendix A). We attribute the Hydro resolution dependence to differences in how $\rm H$}i occupies the largest halos. VN18 found that massive halos in coarser simulations tended to contain less $\rm H$}i than finer ones, in turn suppressing $\rm H$}i clustering. This agrees with other works that studied the resolution dependence of star-formation in IllustrisTNG. For example, M. Donnari et al. (2021) found that the quenching of the massive satellites ($M_{\star}\geq 10^{10}M_{\odot}$) in these massive host halos is enhanced in the coarser IllustrisTNG simulations, in turn suggesting reduced $\rm H$}i abundance.
@@
SAMs exhibit a similar resolution trend in their $\rm H$}i clustering, albeit for different reasons. Instead of a reduction of $\rm H$}i occupation in large halos, SAMs of coarser-resolution simulations possess enhanced $\rm H$}i occupation in small halos (see figure 6 in M. Spinelli et al., 2020). In both SAMs and Hydro, $\rm H$}i occupies smaller halos on average, suppressing its clustering. The $\rm H$}i bias value also depends on the particular SAM implementation, as demonstrated by L. Wolz et al. (2016) and H. Guo et al. (2023). Identifying the particular SAM components that cause these differences is left to future work.
@@
HIHMs, conversely, have a non-trivial dependence on resolution and the employed HIHM. The resolution dependence is clearest when comparing VN18 and Z. Wang et al. (2021), which adopt the same HIHM but still find $\rm H$}i biases that differ by $\sim 0.3$. Z. Wang et al. (2021) reasoned that the discrepancy arises because halos below their resolution threshold contribute non-negligibly to the overall $\rm H$}i distribution. However, even at similar resolutions, D. Sarkar et al. (2016) and Z. Wang et al. (2021) find biases that differ by $\sim 0.5$, implying that the $\rm H$}i bias is also dependent on the adopted HIHM.
@@
Comparisons among the methodologies indicate that $\rm H$}i bias values are sensitive to secondary halo properties, such as halo age or assembly history (L. Gao et al., 2005). Methodologies like Hydro and SAMs that natively incorporate these assembly effects generally converge within $1.2<b_{\mathrm{HI}}<1.5$, particularly at comparable resolutions (R. Ando et al., 2019). A. Marasco et al. (2016) found that SAMs have stronger $\rm H$}i occupation in dense environments as compared to Hydro, which may influence the remaining differences between Hydro and SAMs. Conversely, methodologies that neglect assembly effects are either substantial outliers from Hydro and SAMs, like Analytical, or vary dramatically between implementations, like HIHM. However, in the case of HIHM, VN18 showed that its overall normalization and the fraction of $\rm H$}i outside of halos can also contribute to this disparity in behavior.
@@ 6.2 Why is the FoG effect such a large source of error?
One known auxiliary damping source is the neglected nonlinear Kaiser terms in Equations 12-13, which are typically included as corrections to the matter power spectrum (R. Scoccimarro, 2004; A. Taruya et al., 2010). R. Ando et al. (2019) demonstrated that strictly linear models have reduced PVDs relative to models with these terms (see their figure 5). However, these models necessarily assume that nonlinearities manifest the same in $\rm H$}i and matter (i.e., no $\rm H$}i velocity bias), and furthermore including these nonlinearities can actually worsen model performance at $0.5\lesssim k\lesssim 1$ h cMpc-1. Given these complications, it is uncertain if the nonlinear contributions to the Kaiser term alone can capture the trends seen in Fig. 11.
Differences in the small- and large-scale behavior of the $\rm H$}i PVD may also give rise to the behavior in Fig. 11. The $\rm H$}i PVD can be understood in a halo framework as contributions of three terms, each dominating at different scale regimes: inter-halo (large scales), intra-halo (intermediate), and intra-galactic (small) velocity dispersions (A. Slosar et al., 2006; D. Sarkar & S. Bharadwaj, 2019). The last contribution is unique to $\rm H$}i intensity maps as compared to point-source tracers like galaxies, which would not have intra-galactic dispersions (P23; Z. Li et al., 2024). C. K. Osinga et al. (2024) showed that the $\rm H$}i power spectrum with and without these intra-galactic dispersions diverge at $k\sim 0.5$ h cMpc-1 in IllustrisTNG, with the latter asymptoting to a constant shot noise floor. This comparison demonstrates that intra-galactic dispersions become important at $k\gtrsim 0.5$ h cMpc-1 and the real-space shot noise term is damped only by intra-galactic dispersions [^5] Strictly speaking, shot noise only manifests on scales where random noise emerges and thus cannot be “damped” by RSDs. A more precise interpretation is that $\rm H$}i fills a larger fraction of intra-halo spaces in redshift space, effectively lowering the shot noise floor compared to real space. Consequently, the actual redshift-space shot noise floor should resurface at smaller scales.. Given this conclusion, we propose that $\rm H$}i power spectra be modeled phenomenologically with two separate FoG damping terms:
@@
Here, $P^{\mathrm{SN}}$ symbolizes the shot noise term from the real-space $\rm H$}i auto power spectrum. $\sigma_{\mathrm{HI}}$ represents the $\rm H$}i PVD, which would include all three inter-halo, intra-halo, and intra-galactic components, whereas $\sigma_{\mathrm{int}}$ only includes the intra-galactic dispersions. The two distinct damping terms would reproduce the behavior in Fig. 11 if $\sigma_{\mathrm{int}}<\sigma_{\mathrm{m}}<\sigma_{\mathrm{HI}}$, assuming that shot noise is not significant to the matter power spectrum. A thorough understanding of the distribution of the galactic $\rm H$}i emission profiles can improve the modeling of $\sigma_{\mathrm{int}}$ (Z. Li et al., 2024).
@@ 7 Conclusion
    The error from assuming a constant bias reaches $\gtrsim 10\%$ at $k\sim 0.2-0.3$ h cMpc-1 for $\rm H$}i and all galaxy populations. The exceptions are the $z=1$ $\rm H$}i bias and $z=0$ blue galaxy bias, which remain constant at $10\%$ level until $k\sim 1$ h cMpc-1 because halo occupation effects and nonlinear clustering coincidentally cancel.
@@
    The error from assuming a constant $\rm H$}i-blue galaxy correlation coefficient reaches $10\%$ at $k\sim 0.3$ h cMpc-1. Conversely, $\rm H$}i-red deviates from its large-scale value by $10\%$ on all scales probed by IllustrisTNG ($k\gtrsim 0.05$ h cMpc-1).
@@
However, we only studied the performance of these models using a single suite of simulations. Completing similar studies with other simulations would test how its architecture impacts the resulting tracer distributions and the performance of large-scale $\rm H$}i models. Furthermore, we only examine how models perform in the monopole — higher-order cross-correlations between $\rm H$}i and galaxies can provide even more constraining power, but require even more sophisticated models (e.g., A. Witzemann et al., 2019; E. Chand et al., 2024).
We would like to thank Steven Cunnington and Laura Wolz for their advice, comments, and helpful resources in replicating their work. We also thank Alkistis Pourtsidou and Yi-chao Li for their useful discussions, and the anonymous referee for the helpful comments. Research was performed in part using the compute resources and assistance of the UW-Madison Center For High Throughput Computing (CHTC) in the Department of Computer Sciences. The CHTC is supported by UW-Madison, the Advanced Computing Initiative, the Wisconsin Alumni Research Foundation, the Wisconsin Institutes for Discovery, and the National Science Foundation. We also acknowledge the University of Maryland supercomputing resources (http://hpcc.umd.edu) made available for conducting the research reported in this paper. This research was supported in part by the National Science Foundation under Grant numbers 2206690 and 2338388. BD furthermore acknowledges support from the Sloan Foundation. {contribution}The analysis, writing, and coding was led by CKO. BD wrote the software that created the $\rm H$}i catalogs for IllustrisTNG. FVN provided code that calculated power spectra. Both BD and FVN also advised the project.
@@
The power spectra used in this analysis are provided alongside the online figures: www.calvinosinga.com/papers/hi_cosmo/sup_analysis. Data from specific figures in this publication are available through upon request via email to the authors. The IllustrisTNG data is available at www.tng-project.org/data.